# Quasiparticle cooling algorithms for quantum many-body state preparation

### v1

@@ VI Discussion and outlook
In summary, in this work we investigated cooling protocols for preparing ground states of many-body systems on quantum processors. Our approach is reminiscent of cooling strategies in systems of ultracold atoms, including sympathethic cooling [69], but our protocol primarily aims to remove quasiparticles – emergent excitations above a many-body ground state. We developed a phenomenological kinetic theory framework, which accounts for competing cooling and heating processes induced by auxiliaries and by noise, and compared the efficiency of different cooling protocols.
@@
Our results are directly applicable to studying correlated states with current digital quantum processors, as well as in analog processors with the fast reset capability: indeed, SCP was recently implemented on the Google Quantum Processor [38]. While reaching the ground state in that experiment was challenging due to the level of noise $\gamma\approx 0.01$, our work shows that, with e.g. $\theta=0.1$, noise values of $\gamma\approx 10^{-3}$ will be sufficient for preparing steady states with quasiparticle density below $n\approx 10^{-2}$ in the paramagnetic phase (see Fig. 5b). In systems of $N_{S}=100$ qubits, this translates to approximately one quasiparticle in the sample. Cooling the antiferromagnetic phase is more challenging; nevertheless, noise level $\gamma\approx 10^{-6}$ allows to reach comparable densities. We expect that such low-energy states can be further efficiently purified to extract properties of the ground state and low-energy dynamics [38].
@@
The quasiparticle cooling algorithm is a promising alternative to existing state-preparation methods, including variational unitary circuits [19, 20, 21], thanks to a degree of robustness with respect to noise it offers, as well as its generality. First, dissipation removes excitations created by noise, and the steady state is independent of the initial state and the history of errors. Second, our study of non-integrable TFIM suggests that the cooling approach may be suitable for preparing a broad class of correlated states in non-integrable systems. We note that the cooling algorithm requires the ability to engineer a local coupling operator which has a non-zero overlap with quasiparticle operators. Developing strategies to identify efficient auxiliary-system coupling for models of quantum magnetism, in particular in 2d, is an important direction for future research.
Despite the generality of this approach, we expect certain states will yet prove hard to prepare. One reason is that preparing ground states with high fidelity in general would allow one to solve QMA-complete problems [70], which would contradict a widely believed conjecture that QMA-complete problems cannot be solved efficiently on quantum computers. Still, we do not expect the extreme hard cases to arise in the absence of spin-glass order. In this context, it is therefore important to understand which phases of matter can be prepared by the quasiparticle cooling and other dissipative algorithms.
@@
As pointed out in Ref. [43], and further highlighted by our analysis, systems with topological, non-local quasiparticles are more challenging to cool. On the fundamental level, this stems from lower bounds [71, 72] on the preparation times of topologically ordered states (e.g., toric code). Exploring strategies of optimal cooling of such systems, for example by creating quasiparticle traps or making use of dynamical gauge fields [73], is a promising direction of future research. Furthermore, it would be interesting to apply our analysis to quantum-critical points with gapless quasiparticles, and to study the possibility of cooling strongly interacting systems without long-lived quasiparticles. Generally, we may expect [74] that algorithms similar to ours will be able to efficiently prepare low-energy states of thermalizing systems (which rules out examples such as spin glasses, where cooling procedure will drive the system to one of the many metastable minima).
@@
Our work suggests several other avenues for future research. First, we note that our protocol may be modified to prepare approximate finite-temperature Gibbs states [27, 74, 75, 34], by appropriately choosing the filter function and the operator entering coupling to auxiliary, such that the corresponding rates in Eqs. (42-44) approximately satisfy the detailed balance condition. In fact, this is achieved by our MCP, Eq. (29), see Appendix B, giving a corresponding effective temperature $T_{\rm eff}=1/2\beta$. Weak noise processes in quantum processors generally violate detailed balance, and will make the steady state slightly non-thermal. In contrast, intrinsic thermalization mechanisms in interacting systems are expected to favor the thermal state. This competition, and the resulting steady state may be studied using digital [38] and analog quantum processors.
@@
Finally, our protocol can be modified to engineer non-equilibrium quasiparticle distribution in the steady state. The properties of the steady state, and evolution towards it can give useful insights into interactions between quasipaticles, their dynamics, and lifetimes – properties that play a central role in our understanding of correlated materials. We note that the accuracy of the kinetic theory we used to model the cooling protocol suggests that the dominant thermalization mechanisms are well approximated by few-particle processes. Applications of suitably modified protocols in quantum simulator experiments would unlock new avenues to study underlying dynamics and thermalization mechanisms.
@@ Appendix A Details of modulated coupling pulse
In this Appendix, we provide additional details regarding the modulated coupling protocol (MCP) defined in Eq. (29). As explained in Section II, this pulse is advantageous in that the corresponding filter function, $F_{h,T}(\epsilon)$, approximates a broadened step function, with broadening controlled by the parameter $\beta$. This can be seen as follows: Recalling the definition of the filter function (Eq. 24), we have for the MCP
$F_{h,T}(\epsilon)=\frac{\pi}{A\beta}\sum_{\tau=1}^{T}\frac{\sin[\pi(\tau-\tau_{0})/2]}{\sinh[\pi(\tau-\tau_{0})/\beta]}e^{i\tau(\epsilon-\pi h)}.$ (69)
For the MCP we choose the auxiliary field $h=1/2$. We consider the limits $T,\beta\to\infty$, with the ratio $\alpha=\beta/T\ll 1$ held constant. Then, the summation (69) can be replaced with the integral
[…]
up to an unimportant phase factor. For $\alpha\ll 1$ the integral limits can be extended to infinity and (70) evaluated via residues, giving ($\epsilon>0$)
$F_{h,T}(\epsilon)=\pi\bigg{[}\frac{\tanh(\epsilon\beta/2)-\tanh((\epsilon-\pi)\beta/2)}{\tanh(\pi\beta/4)}\bigg{]}.$ (71)
Here we used the normalization condition $\sum_{\tau=1}^{T}f_{\tau}=1$. The expression for $\epsilon<0$ follows from the residue formula $F(\epsilon)+F(-\epsilon)=2\pi$.
The ratio
$\bigg{|}\frac{F_{h,T}(+\epsilon)}{F_{h,T}(-\epsilon)}\bigg{|}=\bigg{|}\frac{\tanh(\epsilon\beta/2)-\tanh((\epsilon-\pi)\beta/2)}{\tanh(\epsilon\beta/2)-\tanh((\epsilon+\pi)\beta/2)}\bigg{|}$ (72)
$\approx e^{\epsilon\beta},$
gives the detailed balance condition when $\pi\beta\gg{\rm max}(1,\epsilon\beta)$. This condition ensures heating transitions are exponentially less frequent than the cooling transitions of the same energy.
In Fig. 7 we compare $|F_{h,T}(\epsilon)|$ obtained from the integral approximation, (71), and the exact sum (69). We take $\beta=10$ and show the summation result for $T=2\beta$. Even for this relatively small value of $T/\beta$ (or, equivalently, relatively large $\alpha=0.5$), Eq. (71) clearly captures the main behaviour of the filter function. The main discrepancy is due to small oscillations (‘ringing’) on top of the integral result. These finite-time artifacts can be viewed as errors arising from the truncation of the Fourier series limits in Eq. (69) at finite $T$. For large $T/\beta$, this ringing error is bounded approximately as
$|F_{h,T}(\epsilon)-F_{h,\infty}(\epsilon)|\lesssim\frac{\exp(-\pi T/\beta)}{A}.$ (73)
In realistic systems the errors due to noise will typically be much larger than the above truncation errors, and we can obtain efficient cooling with values of $T$, $\beta$, considered in the main text.
Figure 7: Comparison of integral approximation for $F_{h,T}(\epsilon)$ (Eq. 71, black) and exact summation (Eq. 69, orange), for parameters $\beta=10$, $T=2\beta$.
@@ Appendix B Detailed balance and thermal state preparation with MCP
Figure 8: Steady state quasiparticle populations for cooling with thermal MCP, for effective temperature $T_{\rm eff}=1/2\beta$. Results shown are for PM phase with one auxiliary per site, with system size $N_{S}=30$ and reset time $T=5\beta$. The dashed black lines are the expected thermal (Fermi) distributions.
Above, we focused on preparation of ground states, which corresponds to the limit $\beta\Delta\to\infty$, where $\Delta$ is the quasiparticle gap. However, our MCP protocol also provides a way to prepare thermal states (finite temperature Gibbs states) of the same effective Hamiltonian, a subject which has been the focus of several recent studies  [27, 74, 76, 34]. The thermal Gibbs state at an effective temperature $T_{\rm eff}$ is defined by
@@
This can be achieved by cooling the system using the MCP with the parameter $\beta=1/2T_{\rm eff}$, and the additional constraint that the coupling operators $V_{j}$ appearing in Eq. (8) are Hermitian. Then, the general rate equation, Eq. (23), replacing $\hat{\sigma}^{\pm}_{j}$ with the Hermitian operators $V_{j}$, can be written as
@@
$w_{\vec{\alpha}\vec{\beta}}=|F_{h,T}(\Delta(\vec{\alpha},\vec{\beta}))|^{2}|\bra{\vec{\beta}}V_{j}\ket{\vec{\alpha}}|^{2},$ (76)
@@
$\frac{\rho_{\vec{\alpha}\vec{\alpha}}}{\rho_{\vec{\beta}\vec{\beta}}}=\frac{|F_{h,T}(\Delta(\vec{\beta},\vec{\alpha}))|^{2}|\bra{\vec{\alpha}}V_{j}\ket{\vec{\beta}}|^{2}}{|F_{h,T}(\Delta(\vec{\alpha},\vec{\beta}))|^{2}|\bra{\vec{\beta}}V_{j}\ket{\vec{\alpha}}|^{2}},$ (77)
then this provides a solution to the steady state $\delta n_{k}=0$. With the Hermitian constraint and choice $\beta=1/2T_{\rm eff}$, using the asymptotic form of the filter function Eq. (72) gives the required Gibbs state, Eq. (74).
As a demonstration, we calculate the steady state quasiparticle populations from the rate equation in the PM phase of the Floquet TFIM, with one auxiliary per site and keeping single-particle processes as in Section III. We choose hermitian operators $V_{j}=\hat{Z}_{j}$. The results are shown in Fig. 8, for several values of the inverse temperature $\beta$. We take system size $N_{S}=30$ and $T=5\beta$ to reduce ringing artifacts. The results are compared to the expected Fermi distributions which hold for fermionic quasiparticles when Eq. (74) is satisfied; the agreement is generally excellent, suggesting that thermal Gibbs states can be efficiently prepared by the quasiparticle cooling algorithm (where, as in the ground state case, errors due to noise etc. will result in slightly non-equilibrium steady state populations).
@@ Appendix C Quasiparticles in the TFIM
Due to particle-hole symmetry, the eigenvectors of $K_{S}$ with negative quasienergy are related to those of positive quasienergy by the conjugation $\varphi_{k}=\psi_{k}^{*}$, $K_{S}\varphi_{k}=e^{i\epsilon_{k}}\varphi_{k}$. The eigenvectors $\psi_{k}$ and quasienergies $\epsilon_{k}$ can be found by exact diagonalisation of the $K_{S}$ matrix; in Appendix A of Ref. [38], an analytical derivation was also given. Note that the boundary condition defines the quantization condition for the $N_{S}$ quasimomenta $k_{m}$, in terms of which the quasienergies are given by the exact formula
@@ Appendix D Matrix elements for bulk cooling
In Sections III, IV, we considered cooling in the example of the transverse-field Ising model, with two types of matrix elements appearing in the derivation of the rate equations: the $\hat{Z}_{j}$ matrix elements, which we denote as
@@
Although TFIM defines an integrable model, calculation of spin matrix elements can be challenging [77], due to the fact that the Jordan-Wigner transformation between spin and fermion operators is non-local. We will see that this is the case for the $\hat{\sigma}^{\pm}_{j}$ matrix elements below. First, we focus on the simpler case of $\hat{Z}_{j}$:

### v2

@@ VI Discussion and outlook
In summary, in this work we investigated cooling protocols for preparing ground states of many-body systems on quantum processors. Our approach is reminiscent of cooling strategies in systems of ultracold atoms, including sympathetic cooling [78], but our protocol primarily aims to remove quasiparticles – emergent excitations above a many-body ground state. We developed a phenomenological kinetic theory framework, which accounts for competing cooling and heating processes induced by auxiliaries and by noise, discussed how to tailor the protocol parameters to remove quasiparticles from the system, and compared the efficiency of different cooling protocols.
@@
Our results are directly applicable to studying correlated states with current digital quantum processors, as well as in analog processors with the fast reset capability: indeed, SCP was recently implemented on the Google Quantum Processor [38]. While reaching the ground state in that experiment was challenging due to the level of noise $\gamma\approx 0.01$, our work shows that, with e.g. $\theta=0.1$, noise values of $\gamma\approx 10^{-3}$ will be sufficient for preparing steady states with quasiparticle density below $n\approx 10^{-2}$ in the paramagnetic phase (see Fig. 7b). In systems of $N_{S}=100$ qubits, this translates to approximately one quasiparticle in the sample. Cooling the antiferromagnetic phase is more challenging, because the quasiparticles cannot be removed individually by local operators; nevertheless, noise level $\gamma\approx 10^{-6}$ allows to reach comparable densities. We expect that such low-energy states can be further efficiently purified to extract properties of the ground state and low-energy dynamics [38].
@@
The quasiparticle cooling algorithm is a promising alternative to existing state-preparation methods, including variational unitary circuits [19, 20, 21], thanks to a degree of robustness with respect to noise it offers, as well as its generality. First, dissipation removes excitations created by noise, and the steady state is independent of the initial state and the history of errors. Second, our results for the non-integrable models studied here suggest that dissipative cooling should be suitable for preparing a broad class of correlated states, not restricted to integrable systems. Further, we note that the dissipative protocol can be adapted to complement existing state-preparation methods, for example by using the output of variational algorithms for the protocol initial state, or to aid in avoiding local minima in the variational landscape [79].
One area where we expect our cooling protocol to prove instrumental is in the study of 2d correlated systems: the ground states of such systems are challenging to probe numerically, but should be amenable to state preparation on quantum devices. The study of the Heisenberg ladder (a quasi-1d system) carried out in Section IV suggests that the cooling protocol will remain effective also in 2d, in particular for the cases where the system’s excitations can be adiabatically connected to local operators. It will be interesting to understand the advantages or limitations of dissipative cooling for these systems.
@@
As pointed out in Ref. [43], and further highlighted by our analysis, systems with topological, non-local quasiparticles are more challenging to cool. On the fundamental level, this stems from lower bounds [80, 81] on the preparation times of topologically ordered states (e.g., toric code). Exploring strategies of optimal cooling of such systems, for example by creating quasiparticle traps or making use of dynamical gauge fields [82], is a promising direction of future research.
The quasiparticle theory typically breaks down when crossing a quantum phase transition. It would be instructive to apply our analysis to quantum-critical points with gapless quasiparticles, and to study the possibility of cooling strongly interacting systems without long-lived quasiparticles. Although local operators are no longer expected to couple strongly to single-particle excitations, we expect that our cooling protocol should drive the system into a low energy steady state, as long as heating processes are efficiently suppressed (e.g. by the filter function). Generally, we may expect [83] that algorithms similar to ours will be able to efficiently prepare low energy states of thermalizing systems (which rules out examples such as spin glasses, where the cooling procedure will drive the system to one of the many metastable minima). We note that, despite the generality of this approach, we expect certain states will yet prove hard to prepare. One reason is that preparing ground states with high fidelity in general would allow one to solve QMA-complete problems [84], which would contradict a widely believed conjecture that QMA-complete problems cannot be solved efficiently on quantum computers. Still, we do not expect the extreme hard cases to arise in the absence of spin-glass order.
@@
Our work suggests several other avenues for future research. First, we note that our protocol may be modified to prepare approximate finite-temperature Gibbs states [27, 83, 85, 34], by using the MCP filter function, Eq. (38), with effective temperature $T_{\rm eff}=1/\beta$. This is due to the fact that the MCP already satisfies the conditions for detailed balance, Eq. (36). We demonstrate cooling of finite temperature states in Appendix B, leaving a more detailed investigation for future works. Weak noise processes in quantum processors generally violate detailed balance, and will make the steady state slightly non-thermal. In contrast, intrinsic thermalization mechanisms in interacting systems are expected to favor the thermal state. This competition, and the resulting steady state may be studied using digital [38] and analog quantum processors.
@@
Finally, our protocol can be modified to engineer non-equilibrium quasiparticle distribution in the steady state [86]. The properties of the steady state, and evolution towards it can give useful insights into interactions between quasipaticles, their dynamics, and lifetimes – properties that play a central role in our understanding of correlated materials. We note that the accuracy of the kinetic theory we used to model the cooling protocol suggests that the dominant thermalization mechanisms are well approximated by few-particle processes. Applications of suitably modified protocols in quantum simulator experiments would unlock new avenues to study underlying dynamics and thermalization mechanisms.
@@ Appendix A Details of modulated coupling pulse
In this Appendix, we provide additional details regarding the modulated coupling protocol (MCP) defined in Eq. (38). As explained in Section II, this pulse is advantageous in that the corresponding filter function, $F^{\rm MCP}_{h,T}(\epsilon)$, approximately satisfies the thermal detailed balance condition, Eq. (36), with the inverse ‘temperature’ set by the parameter $\beta$. First we derive the form of the MCP, starting from the $2\pi$-periodic function given by
$F_{\beta}(\epsilon)=\frac{\pi}{2A}\sum_{m=-\infty}^{\infty}e^{im\pi}\tanh([\epsilon-m\pi]\beta/4).$ (79)
The summation limit above should be taken as $\lim_{k\to\infty}\sum_{m=-2k}^{2k+1}$, although for $\beta\gtrsim 1$ is is sufficient to restrict to a small number of terms when considering the inverval $-\pi<\epsilon\leq\pi$. By inserting a resolution of the delta-function we arrive to the following Fourier series
$F_{\beta}(\epsilon)=\sum_{m=-\infty}^{\infty}e^{im\epsilon}g(m),$ (80)
$g(m)=\frac{\pi}{2A}\int_{-\infty}^{\infty}\frac{d\epsilon^{\prime}}{2\pi}e^{-im\epsilon^{\prime}}[\tanh(\epsilon^{\prime}\beta/4)-\tanh([\epsilon^{\prime}-\pi]\beta/4)].$
The components can be evaluated via residues,
$g(m)=\frac{\pi}{iA\beta}\frac{(1-e^{-i\pi m})}{\sinh(2\pi m/\beta)},$ (81)
and then, shifting the summation index to $\tau=m+\tau_{0}$ ($\tau_{0}=T/2$) we have (up to an unimportant phase factor)
$F_{\beta}(\epsilon)=\frac{2\pi}{A\beta}\sum_{\tau=-\infty}^{\infty}\frac{\sin[\pi(\tau-\tau_{0})/2]}{\sinh[2\pi(\tau-\tau_{0})/\beta]}e^{i\tau(\epsilon-\pi/2)}.$ (82)
Truncating the series at a finite number of terms $1\leq\tau\leq T$, we arrive at the MCP given in Eq. (38) of the main text. The truncation error is bounded by
$|F_{\beta}-F^{\rm MCP}_{h,T}|\leq 2\sum_{m=\tau_{0}+1}^{\infty}|g(m)|\lesssim A^{-1}e^{-2\pi T/\beta},$ (83)
so is small in the limit $\beta\ll T$. By design, in the prethermal regime $|\epsilon|\ll\pi$ we have
$F_{\beta}(\epsilon)\approx\frac{\pi}{2A}(1+\tanh(\epsilon\beta/4)),$ (84)
which satisfies the detailed balance condition Eq. (36), and therefore so does the MCP filter function, up to the errors given above.
In Fig. 8 we compare $F^{\rm MCP}_{h,T}(\epsilon)$ as obtained from the MCP, with the Fermi function Eq. (79). (For the latter we truncate the series after a few terms with negligible error.) We take $\beta=20$ and show the summation result for $T=\beta$. Even for this large value of $\beta/T$, the truncation (‘ringing’) error is small and $F^{\rm MCP}_{h,T}$ clearly captures the step-like form of the Fermi function. In realistic systems the errors due to noise will typically be larger than the above truncation errors, and we can obtain efficient cooling with values of $T$, $\beta$, considered in the main text.
Figure 8: Comparison of $F^{\rm MCP}_{h,T}$ (Eq. (38), orange) and $F_{\beta}$ (Eq. (79), black), for parameters $T=\beta=20$.
@@ Appendix B Detailed balance and thermal state preparation with MCP
Figure 9: Steady state quasiparticle populations for cooling with thermal MCP, for effective temperature $T_{\rm eff}=1/\beta$. Results shown are for PM phase with one auxiliary per site, with system size $N_{S}=30$ and reset time $T=5\beta/2$. The dashed black lines are the expected thermal (Fermi) distributions.
Above, we focused on preparation of ground states, which corresponds to the limit $\beta\Delta\to\infty$, where $\Delta$ is the excitation gap. However, our MCP protocol also provides a way to prepare thermal states (finite temperature Gibbs states) of the same effective Hamiltonian, a subject which has been the focus of several recent studies  [27, 83, 87, 34]. The thermal Gibbs state at an effective temperature $T_{\rm eff}$ is defined by
@@
This can be achieved by cooling the system using the MCP with the parameter $\beta=1/T_{\rm eff}$, and the additional constraint that the cooling operators $\hat{A}^{a}$ appearing in Eq. (9) are Hermitian. Then, the general rate equation, Eq. (25), with the Hermitian operators $\hat{A}^{a}$, can be written as
@@
$w_{\vec{\alpha}\vec{\beta}}=\sum_{a}|F_{h,T}(\Delta(\vec{\alpha},\vec{\beta}))|^{2}|\bra{\vec{\beta}}\hat{A}^{a}\ket{\vec{\alpha}}|^{2},$ (87)
@@
$\frac{\rho_{\vec{\alpha}\vec{\alpha}}}{\rho_{\vec{\beta}\vec{\beta}}}=\frac{|F_{h,T}(\Delta(\vec{\beta},\vec{\alpha}))|^{2}\sum_{a}|\bra{\vec{\alpha}}\hat{A}^{a}\ket{\vec{\beta}}|^{2}}{|F_{h,T}(\Delta(\vec{\alpha},\vec{\beta}))|^{2}\sum_{b}|\bra{\vec{\beta}}\hat{A}^{b}\ket{\vec{\alpha}}|^{2}},$ (88)
then this provides a solution to the steady state $\delta n_{k}=0$. With the Hermitian constraint and choice $\beta=1/T_{\rm eff}$, using Eq. (36) for the MCP gives the required Gibbs state, Eq. (85).
As a demonstration, we calculate the steady state quasiparticle populations from the rate equation in the PM phase of the Floquet TFIM, with one auxiliary per site and keeping single-particle processes as in Section III. We choose hermitian operators $\hat{A}^{a}=\hat{Z}_{j}$. The results are shown in Fig. 9, for several values of the inverse temperature $\beta$. We take system size $N_{S}=30$ and $T=5\beta/2$ to reduce ringing artifacts. The results are compared to the expected Fermi distributions which hold for fermionic quasiparticles when Eq. (85) is satisfied; the agreement is generally excellent, suggesting that thermal Gibbs states can be efficiently prepared by the quasiparticle cooling algorithm (where, as in the ground state case, errors due to noise etc. will result in slightly non-equilibrium steady state populations).
@@ Appendix C Quasiparticles in the TFIM
Due to particle-hole symmetry, the eigenvectors of $K_{S}$ with negative quasienergy are related to those of positive quasienergy by the conjugation $\varphi_{k}=\psi_{k}^{*}$, $K_{S}\varphi_{k}=e^{i\epsilon_{k}}\varphi_{k}$. The eigenvectors $\psi_{k}$ and quasienergies $\epsilon_{k}$ can be found by exact diagonalisation of the $K_{S}$ matrix; see e.g. Ref. [88] or Appendix A of Ref. [38] for an analytical derivation. Note that the boundary condition defines the quantization condition for the $N_{S}$ quasimomenta $k_{m}$, in terms of which the quasienergies are given by the exact formula
@@ Appendix D Matrix elements for bulk cooling
In Sections III, V, we considered cooling in the example of the transverse-field Ising model, with two types of matrix elements appearing in the derivation of the rate equations: the $\hat{Z}_{j}$ matrix elements, which we denote as
@@
Although the TFIM defines an integrable model, calculation of spin matrix elements can be challenging [89], due to the fact that the Jordan-Wigner transformation between spin and fermion operators is non-local. We will see that this is the case for the $\hat{\sigma}^{\pm}_{j}$ matrix elements below. First, we focus on the simpler case of $\hat{Z}_{j}$: